# A-OctoMap: An Adaptive OctoMap for Online Motion Planning

### v1

@@ III-A Tree build
3:  OctreeNode curNode $\leftarrow$ root
@@
III-B Uniform Grid Mapping
Figure 1: Illustration of mapping from octree to uniform grid. The top row presents the common method of mapping (Fixed Uniform Grid), while the bottom row presents our method (Adaptive Uniform Grid). The red lines represent the split lines of the octree at one level and its parent level. The blue dotted line represents the node boundary of that node concerning the obstacle on it. The red dotted lines represent the split boundary of that node. The right columns show how the uniform grid would appear based on the two methods.
As shown in Fig. 1, we have developed a novel mapping method that does not solely rely on fixed grid sizes. Instead, it determines grid occupancy based on obstacle boundaries. The edge length of the uniform grid is equivalent to the MCR edge length. According to Theorem 1, this boundary must intersect the gap and adjust according to the object boundary.
This adjustment is based on the distance between the SplitBoundary and the NodeBoundary, determining whether a grid should be marked as occupied. In both methods, the edge length of a single uniform grid is equal to half the edge length of the node at the current level. Consequently, this approach does not increase the overall number of grids.
III-C Dynamic Partition
In dynamically changing environments, the limitations of resolution in static uniform grid maps can prevent the guarantee of finding feasible solutions at any given moment. To address this issue, we further subdivide the leaf nodes (see Algorithm 2) in III-C until a feasible solution is identified, and store the obstacles in the specific node leaves.
@@ III-C Dynamic Partition
III-D Downsampling
In our tree-structured architecture, each child node operates independently, enabling parallel computation for recursive operations and synchronous downsampling of each child node. As shown in Algorithm 3 in 2, during the tree construction, the maximum and minimum points in each dimension are determined using the NodeBoundary. These external points are used to construct new convex polyhedra at the leaf nodes. By evaluating the positions of the remaining points relative to each extremal face of convex polyhedra, all points within that are eliminated. Points outside the extremal faces are then treated as new candidate extremal points, facilitating construction using quick hull method.
(a) A 2-D point cloud map built using Perlin Noise
(b) A 3-D point cloud map composed of non-convex geometric shapes
(c) A 3-D point cloud map of a real-world room scan
(d) A 3-D point cloud map of a real-world room scan
Figure 2: 4 scenarios used in our expeirments to present the efficiency of our method
@@ III-D Downsampling
Subsequently, utilizing the Computational Geometry Algorithms Library (CGAL)[21], we apply parallel mesh triangulation to those extreme poinst in each node to represent the map using meshes. These meshes are then stored within their corresponding nodes, enabling efficient and accurate collision detection.
Figure 3: Illustration of mesh triangulation after downsampling.
Our algorithm accesses the gap areas of each node through the node boundaries, using these areas as minimal bounding boxes to avoid maintaining overall density consistency. Compared to 3D Edge-Preserving Sampling, our method retains the extreme points of the convex hull, enabling local mesh triangulation at a lower computational cost while more effectively preserving critical geometric features. Due to the independence of each node, this approach leverages parallel processing to enhance computational efficiency, making it particularly suitable for real-time applications in robotics and autonomous systems.
IV Complexity Analysis and Experiments
IV-A Complexity Analysis
@@
In tree building, the time complexity is the same as the original OctoTree build in average and worst cases. Our tree covers the entire map and maintains balance, preventing points from being inserted into a single path to one leaf node. Both the Uniform Mapping and Dynamic Partition are linear algorithms, dependent on the number of leaf nodes.
For downsampling, we store the point cloud in each corresponding leaf node. The best and average cases of the Quickhull method are linearithmic time, while the worst case is quadratic. In our method, points inside the external face are eliminated, resulting in a computation time of $O\left(\frac{\hat{n}}{m}\log\frac{\hat{n}}{m}\right)$ for each leaf node. The total time complexity is $O\left(\hat{n}\log\frac{\hat{n}}{m}\right)$ in the best and average cases, and $O\left(\frac{\hat{n}^{2}}{m}\right)$ in the worst case.
(a) Original map
(b) VoxelGrid method
(c) Our method
Figure 4: As shown in the three images above, from left to right: the first image is the original point cloud map, the middle image is the point cloud map processed with a voxel filter, and the third image is our method. Our method preserve critical geometric features.
IV-B Experiments
In this section, we present numerical results to demonstrate the effectiveness of our proposed method in four distinct scenarios as shown in Fig. 2, which are based on the RViz in Robot Operation System (ROS) Noetic Ninjemys. We used a Linux desktop with Intel Core i9-13900H running c++ for all computations.
@@ IV-B Experiments
IV-B1 tree build
The table II shows the performance of th tree build in the four scenarios.
@@
IV-B2 uniform mapping and dynamic partition
The table III presents the performance of Uniform Mapping and Dynamic Partition at specific depths of the octo-map. In both Uniform Mapping and Dynamic Partition, the execution times are all in milliseconds, ensuring their viability for online applications.
@@
IV-B3 down sampling and mesh triangulation
The table IV shows the performance of downsampling and mesh triangulation at the 7th depth of the tree. Observing Table IV, we can see that the greater the point cloud size, the lower the retention rate (the retention ratio after downsampling). From Fig. 5, it is evident that both the voxel filter and our method retain 10 % of the points from the original map. However, our method retains only the extremal points of the convex hull for each leaf node, which allows us to better preserve the geometric features of the point cloud while discarding redundant points.
@@
Cell Size (meter) Grid Map Dimension Fixed Uniform Grid Adaptive Uniform Grid (Our method) Planning Time
@@
IV-B4 Jump Point Search
Figure 5 shows that in the same map, JPS generates paths of different lengths in various types of grid maps. Additionally, in Table V, we tested 1000 different maps generated by Perlin noise. From both Figure 5 and Table V, we can observe that our method has a high probability of finding paths in narrow space environments with shorter path lengths. In all 1000 trials, there was no instance where pathfinding succeeded in the uniform grid but failed with our method.
For rigor, we used Eq. (5) to calculate the tree depth, which allowed us to determine the required number of grids and the grid cell size for my method. We then used a uniform grid of the same cell size for comparison, thus eliminating any size-related advantages. We compared the success rates of feasible solution finding and, in cases where paths were found, the path lengths for both methods across different cell sizes. From the Fig. 6, it is evident that in 1000 maps with dimensions of 200 $\times$ 150 meters and approximately 1,200,000 points, there is an overall improvement of at least 6% in the success rate of finding solutions. Additionally, the path length was reduced by more than 6%, and the run time is less than 0.4 ms.
Figure 5: Illustration of JPS in the map, with dimensions of $200\times 150$ meters and approximately 1,200,000 points. The left image (red line path) shows the path finding in the uniform gird map, the right image (blue line path) shows the path finding in the adaptive grid map. The path finding using our method is with shorter path length.
Figure 6: The feasible solution discovery rate and reduction ratio in path length.
@@ V Conclusion and Future Work
This study has presented a comprehensive framework utilizing an innovative OctoMap structure to enhance spatial representation and dynamic partitioning for efficient motion planning in complex environments. Our approach, underpinned by a novel tree-based data structure, offers high-precision segmentation of free spaces and obstacles, facilitating robust and real-time motion planning capabilities. The framework was validated through rigorous experiments in both simulated and real-world scenarios. These experiments demonstrated the system’s capability to navigate cluttered environments and dynamically adjust to new obstacles, outperforming existing methods in terms of speed and reliability. Future efforts will be made to scale the solution to larger environments and more diverse scenarios, including outdoor settings with highly unpredictable elements.

### v2

@@ III-C Tree Construction
3:  OctoTreeNode curNode $\leftarrow$ root
@@
III-D Uniform Grid Mapping
Figure 3: Illustration of mapping from OctoTree to uniform grid. The top row shows the fixed uniform grid method, while the bottom row shows our adaptive method. Red lines indicate OctoTree splits at different levels; red dotted lines represent SplitBoundary, and blue dotted lines represent NodeBoundary. The right columns visualize the resulting grid structures under both methods.
As shown in Fig. 3, we propose a novel mapping method that does not rely solely on fixed-size grids. Instead, grid occupancy is determined by the obstacle’s spatial boundaries, with each grid cell aligned to the SplitBoundary extracted from the tree structure. The edge length of each grid equals that of the MCR, given by $2^{-\text{depth}}\cdot L$, where $L$ is the length of the map domain.
According to Thm. 1, the MCR-based partition ensures that each cell intersects the gap between obstacles, enabling adaptive refinement near object boundaries. Unlike traditional uniform grids that may misalign with obstacle edges, our method guarantees geometric consistency between the grid layout and the map geometry, thus improving representation fidelity without increasing the total number of cells.
A cell is marked as occupied based on the relative distance between its SplitBoundary and NodeBoundary. Since the grid size is derived from the tree depth, it inherits the resolution characteristics of the original node, maintaining efficiency in memory and computation.
III-E Dynamic Partition
We use JPS as the default planner operating on the adaptive grid. During planning, the system interacts with the tree structure to query occupancy; if no feasible path is found, dynamic refinement is triggered. Building on the adaptive mapping framework, we incorporate a dynamic partitioning strategy to handle cases where static uniform grids fail—particularly in environments with emerging obstacles or narrow passages.
When the planner is unable to find a feasible path, the relevant leaf nodes are identified and recursively subdivided to locally increase resolution. New point cloud data can be inserted into the refined nodes, allowing obstacle information to remain up to date. This process continues until a feasible solution is found or the resolution limit is reached.
The dynamic refinement allows the map to adjust to environmental complexity in a localized manner, avoiding the need for full map reconstruction. The recursive update logic is implemented in Algorithm 2.
@@ III-E Dynamic Partition
III-F Downsampling
To reduce point cloud density while preserving geometric features, we propose a structure-aware downsampling method based on local convex hull approximation. Each tree node maintains a local point cloud and operates independently, enabling parallel processing during tree construction.
For each leaf node, we first compute extremal points along each axis to form a convex polyhedron that approximates the surface geometry. Internal points are eliminated, while only representative boundary points are retained to preserve key features. This process is outlined in Algorithm 3.
@@ III-F Downsampling
After filtering, the retained points are used to reconstruct the mesh using the QuickHull algorithm [20] and triangulated via Computational Geometry Algorithms Library (CGAL) [21]. The resulting surface meshes are stored in the respective tree nodes and can be used for downstream planning or visualization.
III-G Complexity Analysis
To better understand the computational efficiency of each module in our proposed framework, we provide a complexity analysis covering the main components: tree construction, uniform (grid) mapping, dynamic partitioning, and downsampling.
Table I summarizes the time complexities in both best-case and worst-case scenarios, with respect to the number of points, tree depth, and dimensionality of the map.
@@ III-G Complexity Analysis
In Tree Construction, the time complexity is the same as the original OctoTree build in average and worst cases. Our tree covers the entire map and maintains balance, preventing points from being inserted into a single path to one leaf node. Both the Uniform (Grid) Mapping and Dynamic Partition are linear algorithms, dependent on the number of leaf nodes.
For Downsampling, we store the point cloud in each corresponding leaf node. The best and average cases of the Quickhull method are linearithmic time, while the worst case is quadratic. In our method, points inside the external face are eliminated, resulting in a computation time of $O\left(\frac{\hat{n}}{m}\log\frac{\hat{n}}{m}\right)$ for each leaf node. The total time complexity is $O\left(\hat{n}\log\frac{\hat{n}}{m}\right)$ in the best and average cases, and $O\left(\frac{\hat{n}^{2}}{m}\right)$ in the worst case. These results highlight that each module of the framework is designed to maintain efficiency even under worst-case conditions, with tree construction and downsampling being the most computationally intensive components due to their recursive and geometry-aware nature.
IV Case Studies
To evaluate the proposed framework under various environmental conditions, we design four representative scenarios involving both synthetic and real-world point cloud maps. These scenarios include 2D Perlin-noise environments, cluttered non-convex geometries, and 3D reconstructions of real indoor spaces. Figure  4 presents the input maps used throughout the following case studies, which are based on the RViz in Robot Operation System (ROS) Noetic Ninjemys. We used a Linux desktop with Intel Core i9-13900H running c++ for all computations. Each of the subsequent subsections focuses on a specific module of our framework, analyzed in the context of these representative environments.
(a) A 2-D point cloud map built using Perlin Noise
(b) A 3-D point cloud map composed of non-convex geometric shapes
(c) A 3-D point cloud map of a real-world room scan (Room 1)
(d) A 3-D point cloud map of a real-world room scan (Room 2)
Figure 4: Four experimental scenarios used to evaluate the effectiveness of the proposed framework.
@@ IV Case Studies
IV-1 Tree Construction
The table II shows the performance of the tree construction in the four scenarios.
@@ IV-1 Tree Construction
IV-2 Uniform (Grid) Mapping and Dynamic Partition
The Table III presents the performance of uniform (grid) mapping and dynamic partition at specific depths of the OctoMap. In both uniform mapping and dynamic partition, the execution times are all in milliseconds, ensuring their viability for online applications.
@@ IV-2 Uniform (Grid) Mapping and Dynamic Partition
IV-3 Downsampling and Mesh Triangulation
We visualize the result of our mesh reconstruction process after downsampling in Fig. 5. The figure shows that our method successfully retains surface geometry while reducing point cloud density. Table IV presents the runtime and retention performance at the $7^{\text{th}}$ depth of the tree. We observe that as the point cloud size increases, the retention rate (i.e., the proportion of retained points) decreases. We also compare our method with the commonly used voxel grid filter [1], which retains a similar percentage of points ($10\%$), as shown in Fig. 6. However, our method selects only the extremal points of the convex hull in each leaf node, enabling better preservation of geometric features while discarding redundant interior points.
Figure 5: Illustration of mesh triangulation after downsampling.
@@ IV-3 Downsampling and Mesh Triangulation
Cell Size (meter) Grid Map Dimension Fixed Uniform Grid Adaptive Mapping Framework (our method) Planning Time
@@
(a) Original map
(b) VoxelGrid method
(c) Our method
Figure 6: As shown in the three subfigures above, from left to right: the left subfigure is the original point cloud map, the middle subfigure is the point cloud map processed with a voxel filter, and the right subfigure is our method. Our method preserve critical geometric features.
IV-4 Jump Point Search
Figure 7 shows that in the same map, JPS generates paths of different lengths in various types of grid maps. Additionally, in Table V, we tested 1000 different maps generated by Perlin noise. From both Fig. 7 and Table V, we can observe that our method has a high probability of finding paths in narrow space environments with shorter path lengths. In all 1000 trials, there was no instance where pathfinding succeeded in the uniform grid but failed with our method.
For rigor, we used Eqn. (2) to calculate the tree depth, which allowed us to determine the required number of grids and the grid cell size for my method. We then used a uniform grid of the same cell size for comparison, thus eliminating any size-related advantages. We compared the success rates of feasible solution finding and, in cases where paths were found, the path lengths for both methods across different cell sizes. From the Fig. 8, it is evident that in 1000 maps with dimensions of 200 $\times$ 150 meters and approximately 1,200,000 points, there is an overall improvement of at least $6\%$ in the success rate of finding solutions. Additionally, the path length was reduced by more than $6\%$, and the run time is less than 0.4 ms.
Figure 7: Illustration of JPS in the map, with dimensions of $200\times 150$ meters and approximately 1,200,000 points. The left subfigure (red line path) shows the path finding in the uniform gird map, the right subfigure (blue line path) shows the path finding in our adaptive grid mapping framework. The path finding using our method is with shorter path length.
Figure 8: Comparison of feasible path discovery rate and path length reduction, showing that our adaptive mapping consistently improves performance, even in cluttered environments.
@@ V Conclusion and Future Work
This paper presents an adaptive grid mapping framework built upon an enhanced OctoMap structure to support efficient and accurate path planning in complex environments. The proposed method introduces a novel tree-based data representation that enables fine-grained segmentation, dynamic spatial refinement, and geometry-aware downsampling. These features collectively facilitate the generation of planning-friendly maps that balance resolution and efficiency. Comprehensive experiments in both simulated and real-world point cloud scenarios demonstrate that our framework achieves higher feasible path discovery rates and produces shorter trajectories compared to traditional uniform-grid-based methods. The framework also supports dynamic updates without full remapping, improving robustness in evolving environments. Future work will explore integration with full planning and control pipelines, scalability to large-scale outdoor scenes, and tighter coupling with real-time sensor data for online operation.